Let V be the vector space of real 2 × 2 matrices with inner product

(A|B) = tr(B

tA).

Let U be the subspace of V consisting of the symmetric matrices. Find an orthog-

onal basis for U

⊥ where U

⊥ = {A ∈ V | (A|B) = 0 ∀B ∈ U}.


1
Expert's answer
2021-02-01T05:42:06-0500

Let

"B=\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}\isin U^\bot"

Then "B" is orthogonal to any element in "U" and, in particular, to the basis elements. Thus,

"Tr(\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}\cdot \begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix})=Tr\begin{pmatrix}
 a & -b \\
 c & -d
\end{pmatrix}=a-d=0"

Thus, "a=d" . Similarly, dotting with the other basis elements, we get the two following equations:

"Tr(\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}\cdot \begin{pmatrix}
 0 & 1 \\
 0 & 0
\end{pmatrix})=Tr\begin{pmatrix}
 0 & a \\
 0 & c
\end{pmatrix}=c=0"

"Tr(\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}\cdot \begin{pmatrix}
 0 & 0 \\
 1 & 0
\end{pmatrix})=Tr\begin{pmatrix}
 b & 0 \\
 d & 0
\end{pmatrix}=b=0"

Thus, using the fact that "a=d" and "c=b=0" , we get that every matrix in "U^\bot" is of the form

"\begin{pmatrix}
 a & 0 \\
 0 & a
\end{pmatrix}"

Thus, "U^\bot" is spanned by the identity matrix.